Pure acid is to be added to a 5% acid solution to obtain 95 liters of 14% solution. What amounts of each should be used?
Answer by Alan3354(69443)

Pure acid is to be added to a 5% acid solution to obtain 95 liters of 14% solution. What amounts of each should be used?
---------
p = pure acid
f = 5%
---
p + f = 95 (total solution)
100p + 5f = 14*95 (total acid)